Which salts will be more soluble in an acidic solution than in pure water? Explain why each salt is or is not more soluble in an acidic solution than in pure water.

CuCl

CsClO4

CaSO4

Be(OH)2

AlPO4
Answer #1
Concepts and reason

The concepts of solubility of a substance and an acidic solution are used in this question. Also, the concepts of an active ion and passive ion are used in this question.

Initially, the salts which are not soluble in an acidic solution than in pure water are determined by using the concepts of active ion, passive ion and solubility of a substance. Then, the salts which will be more soluble in an acidic solution than in pure water are identified among all the given salts.

Fundamentals

The solubility of a substance is defined as the amount of the substance that is dissolved in a particular volume of a solvent. A substance and solvent can be solid, liquid or gas.

An acidic solution is a solution which is acidic in nature, that is, this solution contains more concentration of proton(s) or hydronium ion(s).

Active ion: The ion which will readily react with another oppositely charged ion is called an active ion.

Passive ion: The ion which hindered the reaction with another oppositely charged ion is termed as a passive ion.

The dissociation of copper chloride and cesium chlorate are illustrated as follows.

CuClCu++ClCsClO4Cs++ClO4\begin{array}{l}\\{\rm{CuCl}} \to {\rm{C}}{{\rm{u}}^ + } + {\rm{C}}{{\rm{l}}^ - }\\\\{\rm{CsCl}}{{\rm{O}}_4} \to {\rm{C}}{{\rm{s}}^ + } + {\rm{ClO}}_4^ - \\\end{array}

The copper chloride and cesium chlorate are not more soluble in an acidic solution as they will produce passive anions in an acidic solution. And these passive anions do not readily react with hydronium ions of the acidic solution. Thus, their salts are less soluble in an acidic solution.

The dissociation of calcium sulfate, beryllium hydroxide, and aluminum phosphate are shown below.

CaSO4Ca2++SO42Be(OH)2Be2++2OHAlPO4Al3++PO43\begin{array}{l}\\{\rm{CaS}}{{\rm{O}}_4} \to {\rm{C}}{{\rm{a}}^{2 + }} + {\rm{SO}}_4^{2 - }\\\\{\rm{Be}}{\left( {{\rm{OH}}} \right)_2} \to {\rm{B}}{{\rm{e}}^{2 + }} + 2{\rm{O}}{{\rm{H}}^ - }\\\\{\rm{AlP}}{{\rm{O}}_4} \to {\rm{A}}{{\rm{l}}^{3 + }} + {\rm{PO}}_4^{3 - }\\\end{array}

Since calcium sulfate, beryllium hydroxide and aluminum phosphate will produce active anions in an acidic solution. And these active anions readily react with hydronium ions of the acidic solution. Therefore, their salts are more soluble in an acidic solution.

Ans:

The salts CaSO4,Be(OH)2andAlPO4{\bf{CaS}}{{\bf{O}}_{\bf{4}}}{\bf{, Be}}{\left( {{\bf{OH}}} \right)_{\bf{2}}}{\bf{ and}}{\rm{ }}{\bf{AlP}}{{\bf{O}}_{\bf{4}}} are more soluble in an acidic solution than in pure water.